# D&M: Enriching E-commerce Videos with Sound Effects by Key Moment Detection and SFX Matching

### v1

@@ 4.2 Task-Specific Training
4.2.1 Pre-training for Better Initialization
@@
4.2.2 Tag-aware Negative Sampling
@@
$\text{prob}(s^{-})\propto\left\{\begin{array}[]{cl}sim(u,s^{-}),&\text{if}~{}y(s^{-})\neq y(s),\\ -|sim(u,s)-sim(u,s^{-})|,&\text{if}~{}y(s^{-})=y(s).\end{array}\right.$ (4)
4.2.3 Loss for end-to-end training
@@ 5.1.2 Evaluation Criteria
5.1.3 Implementation Details
All experiments are conducted within the PyTorch framework on eight NVIDIA TESLA V100 32GB GPUs. The size of the video / moment / SFX embeddings is 512. The max length of frames and subtitles is empirically set to 20 and 40, respectively. The number of the learnable queries in DETR-Decoder is 10. We use AdamW with an initial learning rate of 3e-4 with a weight decay of 1e-4. The learning rate is decayed in a cosine schedule strategy. The MSM sub-network for pre-training is trained for 100 epochs at maximum with a mini-batch size of 512. D&M is trained for 300 epochs with a mini-batch size of 256. D&M will be open source.
5.2 Comparison with Baselines
5.2.1 Baselines
  • Two DETR-based methods: Moment-DETR [14] and QD-DETR [22], which can directly localize key moments.
  • Two encoder-only methods: UniVTG [18] and $\text{R}^{2}$-Tuning [20] that predict boundary offsets of each frame and get final key moments through NMS.
[…]
@@ 5.2.1 Baselines
Moment-DETR [14] 31.6 12.0 16.4 3.1 24.0 7.6
QD-DETR [22] 24.1 7.0 7.2 0.6 15.7 3.8
UniVTG [18] 55.8 26.1 37.2 10.8 46.5 18.5
$\text{R}^{2}$-Tuning [20] 55.7 26.5 38.2 11.9 47.0 19.2
@@
5.2.2 Results
@@
5.3 Understanding D&M
Recall our D&M achieves superior performance through proposed moment-SFX matching (MSM) based pre-training and Tag-aware negative sampling (TaNS). We analyse the effectiveness and rationality of these two schemes with several ablation studies. We also analyse the necessity of the individual modalities in the video encoder and the SFX encoder. We report the ${\text{mAP}}_{\text{SFX}}$ ${\text{mAP}}_{\text{vid}}$, and their mean.
5.3.1 The Need of MSM based Pre-training
5.3.2 The Need of TaNS
5.3.3 MSM based Pre-training vs. TaNS
5.3.4 Alternatives to MSM for Pre-training?
We try two alternatives:
$\bullet$ Salience Score Prediction (SSP). SSP is previously used by [14] as an auxiliary task for key moment judgement. In the current context, training SSP allows us to initialize the video encoder.
@@ 5.3.4 Alternatives to MSM for Pre-training?
5.3.5 When to use MSM?
@@
5.3.6 The influence of input modalities for video
@@
5.3.7 The influence of input modalities for SFX
@@ 5.3.8 Qualitative results
6 Summary and Conclusions
We introduce in this paper the new task of video decoration with sound effects (VDSFX) and accordingly develop a task-specific dataset SFX-Moment. To tackle the new task, we propose D&M, an end-to-end network that performs key moment detection and moment-to-SFX matching simultaneously. Extensive experiments on the SFX-Moment dataset allow us to draw conclusions as follows. Given the challenging nature of the task, moment-SFX matching (MSM) based pre-training and tag-aware negative sampling (TaNS) are important for effective training of the D&M network. The proposed method compares favorably against multiple competitive baselines. With the developed dataset and method, our work opens up new possibilities for intelligent E-commerce video analysis and editing.

### v3

@@ 4.2 Task-Specific Training
Pre-training for Better Initialization
@@
Tag-aware Negative Sampling
@@
$\text{p}(s^{-})\propto\left\{\begin{array}[]{cl}sim(u,s^{-}),&\text{if}~{}y(s^{-})\neq y(s),\\ -|sim(u,s)-sim(u,s^{-})|,&\text{if}~{}y(s^{-})=y(s).\end{array}\right.$
Loss for end-to-end training
@@ 5.1 Experimental Setup
Implementation Details
All experiments were conducted using PyTorch on eight NVIDIA TESLA V100 GPUs. The size of the video / moment / SFX embeddings is 256. The max length of frames and subtitles is empirically set to 12 and 40, respectively. The number of the learnable queries in DETR-Decoder is 10. We use AdamW with an initial learning rate of 3e-4 with a weight decay of 1e-4. The learning rate is decayed in a cosine schedule strategy. The MSM sub-network for pre-training is trained with a batch size of 512, while D&M is trained with a batch size of 256.
@@
Moment-DETR 31.6 12.0 16.4 3.1 24.0 7.6
QD-DETR 24.1 7.0 7.2 0.6 15.7 3.8
UniVTG 55.8 26.1 37.2 10.8 46.5 18.5
$\text{R}^{2}$-Tuning 55.7 26.5 38.2 11.9 47.0 19.2
@@
5.2 Comparison with Baselines
Baselines
  • Two DETR-based methods: Moment-DETR (Lei, Berg, and Bansal 2021) and QD-DETR (Moon et al. 2023b), which can directly localize key moments.
  • Two encoder-only methods: UniVTG (Lin et al. 2023b) and $\text{R}^{2}$-Tuning (Liu et al. 2024) that predict boundary offsets of each frame and get final key moments through NMS.
[…]
Results
@@
5.3 Understanding D&M
Recall our D&M achieves superior performance through proposed moment-SFX matching (MSM) based pre-training and Tag-aware negative sampling (TaNS). We analyse the effectiveness and rationality of these two schemes with several ablation studies. We also analyse the necessity of the individual modalities in the video encoder and the SFX encoder. We report ${\text{mAP}}_{\text{SFX}}$, ${\text{mAP}}_{\text{vid}}$, and their mean.
@@ 5.3 Understanding D&M
The Need of MSM based Pre-training
The Need of TaNS
MSM based Pre-training vs. TaNS
Alternatives to MSM for Pre-training?
We try two alternatives:
$\bullet$ Salience Score Prediction (SSP). SSP is previously used by (Lei, Berg, and Bansal 2021) as an auxiliary task for key moment judgement. In the current context, training SSP allows us to initialize the video encoder.
When to use MSM?
@@
The influence of the input modalities for video embedding
@@
The influence of the input modalities for SFX embedding
@@
Efficiency analysis
Given 20 frames and 40 subtitles per video, D&M runs at a throughput rate of 27$\sim$34 videos per second, see Table 7.
GPU Batch size Embedding Decoding Overall Videos per sec.
V100 150 3.34s 1.12s 4.46s 33.6
A10 100 1.88s 1.06s 2.94s 34.0
2080 Ti 50 1.18s 0.67s 1.85s 27.0
Table 7: Inference efficiency.
6 Concluding Remarks
We introduce in this paper a new task of video decoration with sound effects (VDSFX) and accordingly develop a task-specific dataset SFX-Moment. To tackle the new task, we propose D&M, an end-to-end network that performs key moment detection and moment-to-SFX matching simultaneously. Extensive experiments on the SFX-Moment dataset allow us to draw conclusions as follows. Given the challenging nature of the task, moment-SFX matching (MSM) based pre-training and tag-aware negative sampling (TaNS) are important for effective training of the D&M network. The proposed method compares favorably against multiple competitive baselines. With the developed dataset and method, our work opens up new possibilities for intelligent E-commerce video analysis and editing.
Limitations of this study. Currently, each frame is represented by its CLS token. Such a holistic feature might lack the ability to effectively capture small-scale regions of interest. Consequently, for the adaption of D&M to scenarios wherein fine-grained visual details matter, e.g., egocentric or third-person-view instructional videos, the current multi-modal video embedding module has to be improved, say by exploiting patch-level embeddings. Moreover, D&M does not support interactive VDSFX, an important function for its real-world usability. How to enable the model to follow real-time user guidance deserves further investigation.
Acknowledgments
This work was funded by NSFC (No. 62172420), Kuaishou and fund for building world-class universities (disciplines) of Renmin University of China.